OpenStudy (anonymous):

How do I solve using elimination? x^2+y^2=25 x-y=1

OpenStudy (anonymous):

square x= y+1

OpenStudy (anonymous):

that's just the answer.

OpenStudy (jiteshmeghwal9):

Try to make x^2+y^2=(x-y)^2& also try to make no changes to the question\[(x^2+y^2)=(x^2+y^2-2xy)+2xy\]\[\implies x^2+y^2=(x-y)^2+2xy\]now put the value of x -y in the given equation \[\implies x^2+y^2=(1)+2xy\]

OpenStudy (anonymous):

"that's just the answer." ????

OpenStudy (anonymous):

oh wait.. I get it.. sorry.

OpenStudy (jiteshmeghwal9):

now \[1+2xy=25\]\[2xy=25-1\]\[2xy=24\]\[xy=\frac{24}{2}=12\]\[x={12 \over y}\]

OpenStudy (anonymous):

-> x^2 = y^2 +2y+1 and subtract etc...

OpenStudy (jiteshmeghwal9):

so now put the value in the equation\[{12 \over y}-y=1\]now solve this & gt ur answer:)

OpenStudy (anonymous):

thank you much!